OpenStudy (istim):

2 point sources, S1 and S2, oscillating in phase send waves into the air at the same wavelength, 1.98 m. Given that there is a nodal point where the 2 waves overlap, find the smallest corresponding path length difference. Jemurray helped me here already, but I wanna know if a certain formula works too. |PS2−PS1|=(n−1/2)λ/d

OpenStudy (istim):

MY answer was 1m. The correct answer is 0.99m. This is not a rounding problem.

OpenStudy (anonymous):

um isn't the smallest path length difference 0. S1=S2?

OpenStudy (istim):

Sounds agreeable. But then the answer would be less then the correct answer, 0.99.

OpenStudy (anonymous):

like you said, the only cases where two sine waves share common nodal points (zeros) is when they are in perfect phase or perfectly out of phase. so | S1 - S2 | = n λ or | S1 - S2 | = (n-1/2) λ, where the wavelength λ is 1.98. Discarding the trivial case where S1=S2 gives you 1.98/2 = 0.99

OpenStudy (istim):

But I want 0.99....

OpenStudy (istim):

And the formula |PS2−PS1|=(n−1/2)λ/d works in this case right?

OpenStudy (anonymous):

what is d?

OpenStudy (istim):

Distance, like in every other physics equation. Hi Murray. Does the formula |PS2−PS1|=(n−1/2)λ/d work?

OpenStudy (istim):

Cause I used |PS2−PS1|=(n−1/2)λ/d, like the teacher directed, but my answer was off by 0.01.

OpenStudy (anonymous):

There are quite a few distances in the world, so could you be more specific as to what d refers to?

OpenStudy (istim):

I actually dunno. I guess between the 2 points.

OpenStudy (anonymous):

Also I'm not sure I'm doing "path length difference" correctly ... I'm just figuring out the distance between the two point sources = how far they are apart from each other. I'm making them a half-wavelength apart, using (1/2)λ

OpenStudy (istim):

Your guess is better than mine. Hey, Cap Jack! Over here. You're lookin at the wrong question.

OpenStudy (anonymous):

Alright, look, the way this works is the following. You get a dark spot if and only if